OpenStudy (anonymous):

Simplify. ∜400/∜5 A. 5∜2 B. 2∜5 C. 2√5 D. ∜80

OpenStudy (abb0t):

Here's a hint: \[\sqrt[4]{100 \times 4} = \sqrt[4]{(10 \times 10) (2 \times 2)} = \sqrt[4]{(5 \times 2)(5 \times 2)(2 \times 2)}\]

OpenStudy (anonymous):

\[\sqrt[4]{400?}\]

OpenStudy (abb0t):

\[\sqrt[4]{400} = \sqrt[4]{100 \times 4}=...\]

OpenStudy (anonymous):

i'm lost , sorry i don't understand this stuff.

OpenStudy (anonymous):